# Efficient Attention Mechanisms for Large Language Models: A Survey

### v1

@@ 1 Introduction
XAttention [64], SpargeAttn [74] ] ] [ Block-Sparse for Decode [ Quest [54], DoubleSparsity [66], ReSA [53] ] ] [ Routing-based Block Sparsity [ SeerAttention [15], MoBA [33],
InfLLM-v2 [56], NSA [71] ] ] ] [ Clustering Attention [ Attention by Retrieval [ RetrievalAttention [30], ClusterKV [31], MagicPIG [7] ] ] ] [ Others [ Bidirectional Sparse Attention [ BigBird [72], Routing Transformer [47],
@@ 2.2.1 Data-Independent Decay
which shows that contributions from token $n$ are exponentially decayed by the factor $\gamma^{t-n}$ by the time step $t$. Crucially, $\gamma$ is a data-independent decay, which is a fixed parameter for the layer (often one per head in multi-head retention), not a function of the input content. It endows RetNet with an O(1) memory update like an RNN, while still allowing parallel computation during training via an equivalent matrix formulation. (For example, one can show that equation 3 is equivalent to a “retention matrix” form $,\mathrm{Retention}(X)=(QK^{\top}\odot D),V$, where $D_{t,n}=\gamma^{t-n}$ for $t\geq n$ implements the decay and causal masking.)
@@ 2.2.2 Data-Dependent Decay
where $S_{t-1}$ is the previous state, and $G_{t}$ is a gating tensor determined by the token $x_{t}$ If $G_{t}$ is close to 0 in some component, the past state in that component is largely forgotten at time $t$; if $G_{t}\approx 1$, the past is retained. Unlike the constant $\gamma$ in RetNet, here $G_{t}$ varies with $t$ via $x_{t}$. Two notable examples of this strategy in large language model design are Mamba [17, 11] and Gated Linear Attention (GLA) [68].
@@ 3.2.2 Block-Sparse Attention for Decode
Then it selects Top-K blocks with highest $score$ as sparse subset $S$ for attention calculation.
@@ 3.2.3 Routing-based Block-Sparse Attention
MoBA [33] integrate trainable sparse attention into the pretraining stage. It proposes Mixture of Block Attention, which applies the Top-K mechanism from MoE as gating mechanism to decide critical blocks for each query token.
The importance score of each block is computed by the inner product between query token $q$ and the mean pooling result of block $K_{i}$ along the token dimension:
@@ 4.2 Pretrained Models with Hybrid Efficient Attention
With the increasing demand for efficient long-context modeling and diverse computational paradigms, recent research has extensively explored hybrid attention mechanisms. Such strategies combine global and local attention components, often interleaving specialized layers to balance computational cost and performance.
@@
5 Conclusion
This survey presents a comprehensive overview of efficient attention mechanisms, focusing on their algorithmic foundations, practical implementations, and integration into large-scale pre-trained language models. By categorizing linear and sparse attention into well-defined paradigms, we identify key design principles that enable scalability, computational efficiency, and long-context capability. We also analyze how these mechanisms are deployed in state-of-the-art models, either as standalone architectures or as part of hybrid designs that balance local and global computation. As attention-based models continue to scale and diversify, we anticipate further convergence between algorithmic innovation and hardware-aware optimization. We hope this survey provides a useful foundation for future research and system development in the pursuit of efficient, high-performance language modeling.

### v2

@@ 1 Introduction
To structure this survey, we organize the discussion as follows:
  • Section 2 introduces Linear Attention, covering its evolution across different model generations, the associated design principles, and implications for hardware implementation.
  • Section 3 presents Sparse Attention, categorizing sparsity patterns, analyzing deployment scenarios, and offering practical system-level design recommendations.
  • Section 4 reviews Pre-trained Language Models that incorporate efficient attention mechanisms, including both uniform efficient architectures and hybrid models that integrate local, sparse, and dense attention.
  • Section 5 offers an Outlook on future directions, discussing open challenges and potential advances in algorithmic and hardware-aligned research.
@@
XAttention [65], SpargeAttn [75] ] ] [ Block-Sparse for Decode [ Quest [55], DoubleSparsity [67], ReSA [54] ] ] [ Routing-based Block Sparsity [ SeerAttention [15], Landmark [36],
MoBA [33], InfLLM-v2 [57], NSA [72] ] ] ] [ Clustering Attention [ Attention by Retrieval [ RetrievalAttention [30], ClusterKV [31], MagicPIG [7] ] ] ] [ Others [ Bidirectional Sparse Attention [ BigBird [73], Routing Transformer [48],
@@ 2.2.1 Data-Independent Decay
which shows that contributions from token $n$ are exponentially decayed by the factor $\gamma^{t-n}$ by the time step $t$. Crucially, $\gamma$ is a data-independent decay, which is a fixed parameter for the layer (often one per head in multi-head retention), not a function of the input content. It endows RetNet with an O(1) memory update like an RNN, while still allowing parallel computation during training via an equivalent matrix formulation. (For example, one can show that equation 3 is equivalent to a “retention matrix” form $\mathrm{Retention}(X)=(QK^{\top}\odot D),V,$ where $D_{t,n}=\gamma^{t-n}$ for $t\geq n$ implements the decay and causal masking.)
@@ 2.2.2 Data-Dependent Decay
where $S_{t-1}$ is the previous state, and $G_{t}$ is a gating tensor determined by the token $x_{t}$ . If $G_{t}$ is close to 0 in some component, the past state in that component is largely forgotten at time $t$; if $G_{t}\approx 1$, the past is retained. Unlike the constant $\gamma$ in RetNet, here $G_{t}$ varies with $t$ via $x_{t}$. Two notable examples of this strategy in large language model design are Mamba [17, 11] and Gated Linear Attention (GLA) [69].
@@ 3.2.2 Block-Sparse Attention for Decode
Then it selects Top-K blocks with highest scorescore as sparse subset $S$ for attention calculation.
@@ 3.2.3 Routing-based Block-Sparse Attention
Landmark [36] proposes using special landmark tokens to represent each block and trains the attention mechanism to directly retrieve Top-K blocks via these landmark tokens. However, it did not experiment on large-scale pretrained models.
MoBA [33] integrate trainable sparse attention into the pretraining stage. It proposes Mixture of Block Attention, which applies the Top-K mechanism from MoE as gating mechanism to decide critical blocks for each query token. The importance score of each block is computed by the inner product between query token $q$ and the mean pooling result of block $K_{i}$ along the token dimension:
@@ 4.2 Pretrained Models with Hybrid Efficient Attention
With the increasing demand for efficient long-context modeling and diverse computational paradigms, recent research has extensively explored hybrid attention mechanisms. Such strategies combine global and local attention components, often interleaving specialized layers to balance computational cost and performance. The candidate model architecture illustrations are shown in Figure 4.
@@
5 Outlook
This survey presents a comprehensive overview of efficient attention mechanisms, focusing on their algorithmic foundations, practical implementations, and integration into large-scale pre-trained language models. By categorizing linear and sparse attention into well-defined paradigms, we identify key design principles that enable scalability, computational efficiency, and long-context capability. We also analyze how these mechanisms are deployed in state-of-the-art models, either as standalone architectures or as part of hybrid designs that balance local and global computation.
Looking forward, we highlight several key directions that are expected to shape future research in this area:
Architectural Understanding of Hybrid Models
While prior work in linear attention has largely focused on standalone linear architectures, hybrid models are often constructed by combining off-the-shelf linear attention modules with dense or local components. However, it remains unclear whether a stronger linear backbone directly translates to improved hybrid performance. Future work should investigate hybrid models as a distinct architectural class, seeking to understand their composition, interaction effects, and optimization dynamics.
Lossless Sparse Attention and Extended Context
Sparse attention remains challenged by a trade-off between accuracy and computational gain. Fully trained sparse models often underperform dense ones, while post-training sparse approximations face limitations due to lack of end-to-end training. A major research frontier lies in developing sparse attention mechanisms that maintain the expressiveness and accuracy of dense attention, while scaling to much longer contexts. Additionally, the relationship between sparse budget and context length is poorly understood, where fixed top-k schemes may degrade with longer sequences, calling for more adaptive strategies.
Mechanistic Insights into Sparse and Hybrid Attention
While empirical studies have repeatedly demonstrated that hybrid attention models can match or exceed dense models using fewer attention computations, the underlying reasons for this effectiveness remain insufficiently explored. Besides, it is especially important to investigate whether the sparsity patterns that work well in synthetic benchmarks generalize to real-world tasks, and to characterize the limits of sparsity-based generalization.
As attention-based models continue to evolve, we expect further convergence between architectural innovation, theoretical insight, and hardware-aware design. We hope this survey provides a solid foundation for future research into efficient, high-performance language modeling systems.